OpenStudy (anonymous):

suppose you are deposite A dollars in an acount paying an interest rate of r , compounded continuously .Write equation giving the time t needed for your money to double,or doubling time.

OpenStudy (amistre64):

A*e^(rt) = 2A e^(rt) = 2 rt = ln(2) t = ln(2)/r

OpenStudy (anonymous):

ty amistree

OpenStudy (amistre64):

yep :)

OpenStudy (anonymous):

can u have me another post??

OpenStudy (amistre64):

sure

OpenStudy (anonymous):

did u see anothe problem I post another page?

OpenStudy (amistre64):

havent seen it yet

OpenStudy (anonymous):

here : State a rule that could be use to approximate the amout of time t needed to triple the amount of money in a savings account paying r percent interest compounded continuously